OpenStudy (anonymous):

Separating the variables in the equation ye^-x dy/dx = x gives you what equation?

OpenStudy (loser66):

ydy = xe^xdx

OpenStudy (anonymous):

@Loser66 ODE?

OpenStudy (anonymous):

Oh yea, I think so.

OpenStudy (loser66):

hey, if it is ODE, this stuff is the easiest one. good luck, hehehe

OpenStudy (anonymous):

Well.. go figure. I'm terrible at math and made the mistake of joining AP Calc. Lol :(

OpenStudy (loser66):

hahaha. the same with me, I sit on the wrong class so that I am miserable now.

OpenStudy (anonymous):

can you help me with another problem? It's similar to this one but it just goes right through my head. -_- I'll make another post if you want me to.

OpenStudy (loser66):

post

OpenStudy (loser66):

not sure I can help or not,

OpenStudy (anonymous):

Want me to make a new post?

OpenStudy (loser66):

here is ok, as long as I can handle, no need to have a new post. to me, medal is nothing

OpenStudy (anonymous):

Ok one sec, let me type out the equation from my textbook.

OpenStudy (anonymous):

Which of the following is not a differential equation? a. d^2y/dt^2 = t^2 + (y/4) b. 4 + y = 2x + dx/dy c. (x-x^3)dx = 4 - y^2 d. y' = 0 e. t^2 = s + (ds/dt)^2

OpenStudy (loser66):

a,c e

OpenStudy (anonymous):

I think there is only one possible answer.

OpenStudy (loser66):

if so, take off a, and c, just pick e

OpenStudy (anonymous):

Ok, thank you :D

OpenStudy (loser66):

hehehe

OpenStudy (anonymous):

One more question, does dy/dx = dy(x)/dx ? Every time I put dy/dx into wolfram, it keeps showing it as dy(x)/dx.

OpenStudy (loser66):

I don't get the question. when I have to put dy/dx into worlfram, I put y'

OpenStudy (loser66):

and of course, y' w.r.t.x

OpenStudy (anonymous):

Ah, I guess I'll try it that way then. I didn't want to screw up or misplace the variables because sometimes wolfram doesn't understand what I put in at all -_-